OpenStudy (anonymous):

Help please!!! Much appreciated. Write a rational function with vertical asymptotes at x=-2 and x= 1, a horizontal asymptote at y=2, and a zero at 3.

OpenStudy (zpupster):

vertical asymptoes for x must be in the denominator: f(x) = g(x)/ (x+2)(x-1) g(x) must contain the same degree as the denominator becasue f has a horizantal asymptotes. so 3(x-2)/ (x+2)(x-1)

OpenStudy (anonymous):

Is that the answer then?

OpenStudy (anonymous):

no that one has a zero at 2, you want a zero at 3

OpenStudy (anonymous):

denominator should be \((x+2)(x-1)\) as stated above

OpenStudy (anonymous):

since that is a polynomial of degree 2, your numerator has to be a polynomial of degree 2 as well, and the leading coefficient has to be 2 to give you the correct horizontal asymptote

OpenStudy (anonymous):

first guess would be \[\frac{2x^2}{(x+2)(x-1)}\] but unfortunately that has a zero at \(x=0\) and you need one at \(x=2\)

OpenStudy (anonymous):

try \[\frac{2(x-2)^2}{(x+2)(x-1)}\]

OpenStudy (anonymous):

Then what @satellite73 ?